write in equation in vertex form and identify the conic section and graph it.
x^2+y^2+14y=-13

Answer by stanbon(57395)

write in equation in vertex form and identify the conic section and graph it.
x^2+y^2+14y=-13
-----
x^2 + (y^2+14y+49) = -13+49
----
(x-0)^2 + (y+7)^2 = 36
---
Draw a circle with center at (0,-7) and radius = sqrt(36) = 6
=============
cheers,
Stan H.
==================